OpenStudy (anonymous):

how to multiply 8+4sqrt(2)-4sqrt(7) with 4sqrt(2)+1

OpenStudy (anonymous):

\[8+4\sqrt{2}-4\sqrt{7}\times4\sqrt{2}+1\]

OpenStudy (anonymous):

I use mathematica (wolfram alpha) software for equational and graphical fit modeling for statistics and regressions and to check if my answers and of course to have neat graphs... but I wouldn't substitute it for really learning how to solve simple equations

OpenStudy (id21):

\[(8+4\sqrt{2}-4\sqrt{7})(4\sqrt{2}+1)=8\times4\sqrt{2}+4\sqrt{2}\times4\sqrt{2}-4\sqrt{7}\times 4\sqrt{2}+\\+8\times1+4\sqrt{2}\times1-4\sqrt{7}\times1=\\=32\sqrt{2}+32-16\sqrt{7\times2}+8+4\sqrt{2}-4\sqrt{7}=\\=36\sqrt{2}-16\sqrt{14}-4\sqrt{7}+40\]

OpenStudy (unklerhaukus):

id21 has the solution,\[\color{red}\checkmark\] if your going to use wolfram to check , you need to make sure you enter the right question http://www.wolframalpha.com/input/?i=%288%2B4sqrt%282%29-4sqrt%287%29%29+with+%284sqrt%282%29%2B1%29